what is the error propogation for z = arctan((y/2)/L)

and what would the error propogation be for (d * sin(z)/m)

where

y = .51mm +-.5mm

L = 2m +- .2mm

d = .25 mm

and m = 5

please show steps
Answer #1

We have, Y = 0.51 mm & 0.5 mm So, ay 6.5 mm 2 m 0.2 mm + mm = 2000 mm so, CL a 0.2 mm tant (2/2) and, Z = 52? У d2 ={1331 °+{

6.24 x 10-6 So, Sz = 1. 25x10 -4 Now let us Concider Sinz dx X = Х m ma 5 and Here d = 0-25 mm SZ Cos Z So, Sx a VE tanrı (7)